trimethaphan is administered in the form of trimethaphan camsylate for the rapid reduction of blood pressure. an iv infusion of trimethaphan is administered to a child weighing 75 pounds at a rate of 0.1 mg/min per kilogram if the childs mass. howmmany milligrams of trimethaphan are administered during a one hour infusion.
Answer #1

ANSWER.

* Calculation of mg of Trimethaphan administered during 1 hour infusion.

              Ordered dose = 0.1 mg/kg/min.

              Weight of the child = 75 pounds.

   Convert pound to kg for calculation

            1 pound = 0.454 kg.

So   

         Weight of the child = 75 * 0.454

                                        = 34.05 kg.

     Dose required for a child with 75 pound weight = 0.1 * 34.05

                                                                   = 3.4 mg/min.

     Dose administered during 1 hour infusion = 3.4 * 60 min.

                                                                  = 204 mg.